A ball of mass 0.2kg falls from a height of 45m. On striking the ground, it rebounds in 0.1s with ⅔ of the velocity with which it struck the ground. Calculate (i) the momentum change on hitting the ground (ii) the force on the ball due to the impact
1
Expert's answer
2020-11-18T09:25:48-0500

Applying the principle of conservation of energy,

The potential energy of the ball before falling is equal to the kinetic energy of the ball right before it strucks the ground.


"mgh=\frac{1}2m{v_1}²\\
v_1=\sqrt{2gh}\\"


"\begin{aligned} \textsf{where }&m=\textsf{mass of the ball}\\ &g=\textsf{acceleration due to gravity}\\&h=\textsf{height from which the ball fell}\\&v_1=\textsf{velocity with which the ball struck the ground}\\\end{aligned}" "v_1=\sqrt{2gh}\\
v_1=\sqrt{2×9.81×45}\\
v_1=9.4ms^{-1}\\
\textsf{Considering the direction,} \overrightarrow{v_1} = -9.4ms^{-1}.\\
\textsf{Let } v_2 \textsf{represent the velocity with which the ball rebounds}\\
\overrightarrow{v_2}=-\frac{2}3\overrightarrow{v_1}\\
\overrightarrow{v_2}=-\frac{2}3(-9.4ms^{-1})\\
\overrightarrow{v_2}=6.26ms^{-1}\\"

(i) the momentum change on hitting the ground is,

"\begin{aligned} \Delta P&=m\overrightarrow{v_2}-m\overrightarrow{v_1}\\&=m(\overrightarrow{v_2}-\overrightarrow{v_1})\\&=0.2kg(6.26ms^{-1}-(-9.4ms^{-1}))\\\Delta P&=3.132Ns\end{aligned}"

(ii) The change of momentum is equal to the impulse. The impulse is given by,

"I=Ft\\
\begin{aligned}\textsf{Where }&F=\textsf{force due to impact}\\&t=\textsf{time of impact} \end{aligned}\\\hspace{1cm}\\Ft=\Delta P\\
F=\frac{\Delta P}t\\
F=\frac{3.132Ns}{0.1s}\\
F=31.32N"